# Doping $S=1$ antiferromagnet in one-dimension

### v1

@@ Abstract
Antiferromagnetic ground states, when doped, give rise to rich and complex phenomena, prompting detailed investigations in various spin systems. Here, we study the effect of doping on the one-dimensional $S=1$ antiferromagnetic Heisenberg model (AFM). Specifically, we investigate how the presence of holes affects the static and dynamic (frequency-dependent) spin-spin correlations of the two-orbital Hubbard-Kanamori chain. The latter, at half-filling and in the limit of strong interactions, maps onto an $S=1$ Heisenberg model. For moderate interactions, an orbital resonating-valence-bond (orbital-RVB) state emerges up to doping levels of $x\lesssim 0.4$. A detailed analysis of interaction strength $U$ and doping concentration $x$ reveals that this phase inherits the key features of spin excitations found in the half-filled case — namely, a gapped spin spectrum and “coherent” magnon behavior up to a wavevector $q$ determined by the Fermi vector, $2k_{\mathrm{F}}=\pi(1-x)$. Furthermore, our results uncover an additional broad, incoherent spectral weight for $q\gtrsim 2k_{\mathrm{F}}$ at high frequencies. Finally, we show that near the transition to a ferromagnetic phase, a previously unidentified spiral-like state emerges, characterized by spin excitations reminiscent of the $J_{1}$-$J_{2}$ Heisenberg model.
@@ I Introduction
In a 1D AFM system of $S=1/2$ fermions, introducing charge carriers can significantly alter the magnetic properties, giving rise to incommensurate spin correlations [31], spin-charge separation [32, 11], and even superconducting tendencies under certain conditions. Understanding how doping affects $S=1$ AFM chains is especially intriguing due to the presence of the Haldane gap and the topological nature of the undoped ground state [33, 34, 35]. Canonically, the ground-state of the $S=1$ Heisenberg model is explained via the connection to the Affleck-Kennedy-Lieb-Tasaki (AKLT) valence bond (VB) state with only nearest-neighbor spin-spin correlations [36]. The isotropic $S=1$ Heisenberg model is a gapped state with exponentially decaying correlations without any spin long-range order [the state akin to resonating-VB (RVB)]. In principle, introducing holes (or electrons) into RVB encapsulates Phil Anderson’s idea [37, 38] for the formation of Cooper pairs and eventually for the superconducting state in cuprates [39].
It was recently shown [40] that the RVB-nature of $S=1$ ground state survives till quite large doping at interaction strength comparable to the kinetic energy and, indeed, leads to the superconducting tendencies [41, 42, 18]. Such a phase (at finite doping) was coined the orbital-RVB state. In this work, we will focus on a detailed analysis of the magnetic properties of the latter state. We achieved this by investigating the static and dynamic (frequency-dependent) spin-spin correlations within the two-orbital Hubbard-Kanamori model. The Schrieffer-Wolff transformation of the latter in the large interaction limit (and at half-filling) results in an $S=1$ Heisenberg model. Our findings indicate that spin excitations crucially depend on the interaction strength and doping $x$. Within the orbital-RVB phase, the low-frequency spin spectrum resembles that of an $S=1$ antiferromagnet (with ”coherent” magnon mode and a spin gap). However, at large frequencies, an additional feature appears as a broad, incoherent spectrum at short wavelengths. Finally, we also discuss the spiral-like state that emerges at the transition between orbital-RVB and ferromagnetic phases.
The paper is structured as follows. In Section II, we introduce the two-orbital Hubbard-Kanamori model together with the main quantities and the methods used in this work. In Section III, we describe how the $S=1$ ground state emerges in the two-orbital model and discuss the general properties of the doped $S=1$ state. Section IV contains the main results of our work, i.e., the analysis of the static (in Sec. IV.1) and dynamic (in Sec. IV.2) spin structure factors. The conclusions are given in Section V.
@@ II Model & Methods
In this work, we will focus on the two-orbital Hubbard-Kanamori (HK) model in one spatial dimension:
@@
Here $c^{\dagger}_{\gamma\ell\sigma}$ ($c^{\phantom{\dagger}}_{\gamma\ell\sigma}$) represent electron creation (annihilation) operator at orbital $\gamma=\{0,1\}$, site $\ell=\{1,\dots,L\}$, and with spin projection $\sigma=\{\uparrow,\downarrow\}$. $\Delta_{\mathrm{CF}}$ stands for crystal field splitting, $n_{\gamma\ell}=n_{\gamma\ell\uparrow}+n_{\gamma\ell\downarrow}$ is density of the electrons, $U$ ($U^{\prime}=U-5/2J_{\mathrm{H}}$) is intra-orbital (inter-orbital) Hubbard interaction, $J_{\mathrm{H}}$ is the Hund exchange coupling, $\mathbf{s}_{\gamma\ell}=(1/2)\sum_{a,b}c^{\dagger}_{\gamma\ell a}\mathbf{\sigma}_{ab}\,c^{\phantom{\dagger}}_{\gamma\ell b}$ is the spin operator (with $\mathbf{\sigma}_{ab}$ as the Pauli spin matrices), and $P^{\dagger}_{\gamma\ell}=c^{\dagger}_{\gamma\ell\uparrow}c^{\dagger}_{\gamma\ell\downarrow}$ represent pair hopping operator. In the following, we will use a noninteracting kinetic energy span $W=4t$ as a unit of energy. We will consider an open-boundary system at fixed magnetization $S^{z}_{\mathrm{tot}}=\sum_{\ell,\gamma}s_{\gamma,\ell}=0$ and various hole doping levels x=1nx=1-n, where $n=\sum_{\gamma\ell}n_{\gamma\ell}/L=N/2L$ (with $N$ as the total number of electrons and $L$ as the system size). Since both Hubbard and Hund interactions originate from Coulomb many-body forces, we will fix the Hund coupling to $J_{\mathrm{H}}/U=0.25$ throughout the work. Past analysis of the HK model suggests that varying the $J_{\mathrm{H}}/U$ ratio does not significantly alter the conclusions [42, 43].
@@ III Doping $S=1$ antiferromagnet
The studies of (1) model indicated [43] that such $S=1$ description is valid not only in the $U\to\infty$ limit, but already for rather low values of interaction $U>U_{\mathrm{c}}\simeq W/2$. The latter marks the interaction-induced phase transition between trivial and topological (akin to the AKLT state) ground-states. Such a transition is manifested by the appearance of, e.g., spin gap $\Delta_{\mathrm{S}}\neq 0$, edge states, and finite string order parameter. On the other hand, for finite doping $x\neq 0$ and at $U,J_{\mathrm{H}}\gg t$, the double-exchange mechanism dominates and the system orders ferromagnetically [50, 51, 52]. The analysis of the spin excitations in this region was recently shown in Ref. [53]. See Fig. 1 for sketch of the interaction-doping, $U$-$x$, phase diagram.
@@
It is worth noting that lifting band degeneracy (e.g., by $\Delta_{\mathrm{CF}}\neq 0$ or $t_{11}\neq t_{00}$) may result in other exotic phases. In the case of orbital differentiation, one of the orbitals can undergo a Mott phase transition. At the same time, the other remains delocalized (metallic), i.e., the system can undergo a so-called orbitally-selective Mott transition (OSMP) [54, 25]. Here, the competition between superexchange (AFM) and double-exchange (FM) mechanisms leads to block [55, 56, 57, 58, 59, 60] and block-spiral [61] magnetic order. As we will show below, the remnants of some of these phases are also present for degenerated bands.
@@
Note that the HK model (1) is particle-hole symmetric, i.e., our results are valid also for electron doping, $n=1+x$ (provided that one considers degenerated bands). The above states differ from those expected in the ”standard” single-band Hubbard model. Here, each holon (the site where one of the orbitals is empty) carries a $\uparrow,\downarrow$-spin flavor. Consequently, while the doping of the single-orbital Hubbard model introduces spinless holons (or doublons) in the $S=1/2$ AFM background (encapsulated in the $t$-$J$ model), for the two-orbital Hubbard model, doping introduces $S=1/2$-like objects in the $S=1$ AFM background. Note also that the atomic level allows for an empty site (both orbitals empty), provided that the overall desired $n$ is achieved. However, our analysis indicates that such states have vanishing weight in the ground state. In Fig. 2, we show the doping dependence of the density of various atomic-level states. We use the standard definition of (here orbital resolved) singlons and holons, i.e., $n_{\gamma\ell}^{\mathrm{S}}=n_{\gamma\ell\uparrow}+n_{\gamma\ell\downarrow}-2n_{\gamma\ell\uparrow}n_{\gamma\ell\downarrow}$ and $n_{\gamma\ell}^{\mathrm{H}}=1-n_{\gamma\ell\uparrow}-n_{\gamma\ell\downarrow}+n_{\gamma\ell\uparrow}n_{\gamma\ell\downarrow}$, respectively. Next, we define density of site singlons $n_{\mathrm{SS}}=\sum_{\ell}n_{0\ell}^{\mathrm{S}}\,n_{1\ell}^{\mathrm{S}}/L$ (sites occupied by two electrons, each at one of the orbitals), orbital singlons $n_{\mathrm{SH}}=\sum_{\ell}\left(n_{0\ell}^{\mathrm{S}}\,n_{1\ell}^{\mathrm{H}}+n_{1\ell}^{\mathrm{S}}\,n_{0\ell}^{\mathrm{H}}\right)/L$ (sites occupied by one electron at one of the orbitals), and site holons $n_{\mathrm{HH}}=\sum_{\ell}n_{0\ell}^{\mathrm{H}}\,n_{1\ell}^{\mathrm{H}}/L$. It is evident from the presented results that in the $U\gg t$ limit only site singlons and orbital singlons survive.
@@ IV Magnetic properties
As discussed in the previous section, it was shown [43, 40] that the above $S=1$ Heisenberg model description (with characteristic for such system finite spin gap $\Delta_{S}$, topologically protected edge states, and string order parameter) is valid not only in the $U\gg t$ limit, but also in the regions where the magnetic moments are not yet fully developed $\mathbf{S}^{2}=S(S+1)\neq 2$. For $x=0$ this is true for $U>U_{\mathrm{c}}$, while for $x\neq 0$ in the so-called orbital-RVB state, i.e., for $x\lesssim 0.3$ and $U\sim{\cal O}(W)$. Interestingly, our analysis of the average magnetic moment, $\mathbf{S}^{2}=(1/L)\sum_{i}\mathbf{S}^{2}_{i}$, indicates that in the orbital-RVB state the magnetic moments are never maximized to the $\mathbf{S}^{2}_{\mathrm{max}}(x)=2\cdot(1-2x)+0.75\cdot 2x$ value. The data presented in Fig. 3 indicates that $\mathbf{S}^{2}=\mathbf{S}^{2}_{\mathrm{max}}(x)$ only for $x=0\,,U\gg t$ and when the system orders ferromagnetically for $x\neq 0$.
@@ IV.1 Static structure factor $S(q)$
In contrast to $x=0$ - with one sharp phase transition between the trivial (paramagnetic) state and topological AKLT-like state at $U_{\mathrm{c}}$ - the finite doping $x\neq 0$ exhibits two transitions [40]. The lower one starts at $U_{\mathrm{c}}=W/2$ at $x=0$ and slowly increases till $U_{\mathrm{c}}(x)\sim 4W$ at $x\simeq 0.35$. At large $U$ (and consequently large $J_{\mathrm{H}}$), the system undergoes a second, upper transition, entering a ferromagnetic state with trivial topological properties. Our analysis of static SSF indicates that the latter transition is marked by a continuous change of the position maximum of $S(q)$ from $q_{\mathrm{max}}=2k_{\mathrm{F}}$ (incommensurate AFM) to $q_{\mathrm{max}}=0$ (ferromagnet) [see Fig. 4(b) and Fig. 4(d)]. Such behavior is akin to the spiral phase (with $\langle\mathbf{S}_{i}\times\mathbf{S}_{i+1}\rangle\neq 0$ instability) found in systems exhibiting OSMP, i.e., with lifted band degeneracy [61]. Here, for example, for $x=0.1$, we find a wide range of interaction, $3\lesssim U\lesssim 10$, where such a state is stabilized. Note that the range of interaction values that support the spiral shrinks as the doping $x$ increases [see Fig. 4(d) and sketch in Fig. 1].
@@ IV.2.1 Doping dependence
Such vanishing weight of short wavelengths $q\to\pi$ spectrum and high-$\omega$ modes were already reported in the case of the spin excitations of the OSMP systems [58, 60] (i.e., with orbital differentiation). Also, experimental inelastic neutron scattering (INS) investigation of the low-dimensional OSMP compound BaFe2Se3 [57] indicates on the presence of such modes. There, the high-frequency modes take the form of well-defined (”coherent”), momentum-independent excitations. It was shown [60] that the latter depend on the value of the Hund exchange $J_{\mathrm{H}}$. Here, in the case of the doped $S=1$ Heisenberg model, the high-frequency mode is incoherent and dispersive. Nevertheless, the large frequency span $\sim 20J\sim J_{\mathrm{H}}$ could indicate that the Hund exchange also controls these high-$\omega$ features of the spectra.
IV.2.2 Interaction dependence

### v2

@@ Abstract
Antiferromagnetic ground states, when doped, give rise to rich and complex phenomena, prompting detailed investigations in various spin systems. Here, we study the effect of doping on the one-dimensional $S=1$ antiferromagnetic Heisenberg model (AFM). Specifically, we investigate how the presence of holes affects the static and dynamic (frequency-dependent) spin-spin correlations of the two-orbital Hubbard-Kanamori chain. The latter, at half-filling and in the strong-interaction limit, maps onto an $S=1$ Heisenberg model. For moderate interactions, an orbital resonating-valence-bond (orbital-RVB) state emerges up to doping levels of $x\lesssim 0.3$. A detailed analysis of interaction strength $U$ and doping concentration $x$ reveals that this phase inherits the key features of spin excitations found in the half-filled case — namely, a gapped spin spectrum and “coherent” magnon behavior up to a wavevector $q$ determined by the Fermi vector, $2k_{\mathrm{F}}=\pi(1-x)$. Furthermore, our results uncover an additional broad, incoherent spectral weight for $q\gtrsim 2k_{\mathrm{F}}$ at high frequencies. Finally, we show that near the transition to a ferromagnetic phase, a previously unidentified spiral-like state emerges, characterized by spin excitations reminiscent of the $J_{1}$-$J_{2}$ Heisenberg model.
@@ I Introduction
In a 1D AFM system of $S=1/2$ fermions, introducing charge carriers can significantly alter the magnetic properties, giving rise to incommensurate spin correlations [31], spin-charge separation [32, 11], and even superconducting tendencies under certain conditions. Understanding how doping affects $S=1$ AFM chains is especially intriguing due to the presence of the Haldane gap and the topological nature of the undoped ground state [33, 34, 35]. Canonically, the ground-state of the $S=1$ Heisenberg model is explained via the connection to the Affleck-Kennedy-Lieb-Tasaki (AKLT) valence bond (VB) state with only nearest-neighbor spin-spin correlations [36]. The isotropic $S=1$ Heisenberg model is a gapped state with exponentially decaying correlations without any spin long-range order [akin to the resonating valence bond (RVB) state]. In principle, introducing holes (or electrons) into the RVB state encapsulates Phil Anderson’s idea [37, 38] for the formation of Cooper pairs and, eventually, the superconducting state in cuprates [39].
It was recently shown [40] that the RVB-nature of $S=1$ ground state survives till quite large doping at interaction strength comparable to the kinetic energy and, indeed, leads to the superconducting tendencies [41, 42, 18]. Such a phase (at finite doping) was coined the orbital-RVB state. In this work, we will focus on a detailed analysis of the magnetic properties of the latter state. We achieved this by investigating the static and dynamic (frequency-dependent) spin-spin correlations (i.e., the dynamical spin structure factor) within the two-orbital Hubbard-Kanamori model. The Schrieffer-Wolff transformation of the latter in the large interaction limit (and at half-filling) yields an $S=1$ Heisenberg model. Our findings indicate that the spin excitations crucially depend on the interaction strength and doping $x$. Within the orbital-RVB phase, the low-frequency spin spectrum resembles that of an $S=1$ antiferromagnet (with ”coherent” magnon mode and a spin gap). However, at large frequencies, an additional feature appears as a broad, incoherent spectrum at short wavelengths. Finally, we also discuss the spiral-like state that emerges at the transition between orbital-RVB and ferromagnetic phases.
The paper is structured as follows. In Section II, we introduce the two-orbital Hubbard-Kanamori model together with the main quantities and the methods used in this work. In Section III, we describe how the $S=1$ ground state emerges in the two-orbital model and discuss the general properties of the doped $S=1$ state. Section IV presents the main results of our work: the analysis of the static (Sec. IV.1) and dynamic (Sec. IV.2) spin structure factors. The conclusions are given in Section V.
@@ II Model & Methods
In this work, we will focus on the two-orbital Hubbard-Kanamori (HK) model [43, 25, 44] in one spatial dimension:
@@
Here $c^{\dagger}_{\gamma\ell\sigma}$ ($c^{\phantom{\dagger}}_{\gamma\ell\sigma}$) represent electron creation (annihilation) operator at orbital $\gamma=\{0,1\}$, site $\ell=\{1,\dots,L\}$, and with spin projection $\sigma=\{\uparrow,\downarrow\}$. $\Delta_{\mathrm{CF}}$ stands for crystal field splitting, $n_{\gamma\ell}=n_{\gamma\ell\uparrow}+n_{\gamma\ell\downarrow}$ is density of the electrons, $U$ ($U^{\prime}=U-5/2J_{\mathrm{H}}$) is intra-orbital (inter-orbital) Hubbard interaction, $J_{\mathrm{H}}$ is the Hund exchange coupling, $\mathbf{s}_{\gamma\ell}=(1/2)\sum_{a,b}c^{\dagger}_{\gamma\ell a}\mathbf{\sigma}_{ab}\,c^{\phantom{\dagger}}_{\gamma\ell b}$ is the spin operator (with $\mathbf{\sigma}_{ab}$ as the Pauli spin matrices), and $P^{\dagger}_{\gamma\ell}=c^{\dagger}_{\gamma\ell\uparrow}c^{\dagger}_{\gamma\ell\downarrow}$ represent pair hopping operator. In the following, we will use a noninteracting kinetic energy span $W=4t$ as a unit of energy. We will consider an open-boundary system at fixed magnetization $S^{z}_{\mathrm{tot}}=\sum_{\ell,\gamma}s_{\gamma,\ell}=0$ and various hole doping levels x=1nx=1-n, where $n=\sum_{\gamma\ell}n_{\gamma\ell}/L=N/2L$ (with $N$ as the total number of electrons and $L$ as the system size). Since both Hubbard and Hund interactions originate from Coulomb many-body forces, we will fix the Hund coupling to $J_{\mathrm{H}}/U=0.25$ throughout the work. Past analysis of the HK model suggests that varying the $J_{\mathrm{H}}/U$ ratio does not significantly alter the conclusions [42, 45]. Nevertheless, for completeness, at the end of Sec. IV.2 we present results for various $J_{\mathrm{H}}/U$ ratio.
@@ III Doping $S=1$ antiferromagnet
The studies of (1) indicated [45] that such $S=1$ description is valid not only in the $U\to\infty$ limit, but already for rather low values of interaction $U>U_{\mathrm{c}}\simeq W/2$. The latter marks the interaction-induced phase transition between trivial and topological (akin to the AKLT state) ground-states. The transition is manifested by the appearance of, e.g., spin gap $\Delta_{\mathrm{S}}\neq 0$, edge states, and finite string order parameter. On the other hand, for finite doping $x\neq 0$ and at $U,J_{\mathrm{H}}\gg t$, the double-exchange mechanism dominates and the system orders ferromagnetically [52, 53, 54]. The analysis of the spin excitations in this region was recently shown in Ref. [55]. See Fig. 1 for sketch of the interaction-doping, $U$-$x$, phase diagram.
@@
It is worth noting that lifting band degeneracy (e.g., by $\Delta_{\mathrm{CF}}\neq 0$ or $t_{11}\neq t_{00}$) may result in other exotic phases. In the case of orbital differentiation, one of the orbitals can undergo a Mott phase transition. At the same time, the other remains delocalized (metallic), i.e., the system can undergo a so-called orbitally-selective Mott transition (OSMP) [56, 25]. Here, the competition between superexchange (AFM) and double-exchange (FM) mechanisms leads to block [57, 58, 59, 60, 61, 62] and block-spiral [63] magnetic order. As we show below, remnants of some of these phases are also present in degenerated bands.
@@
Note that the HK model (1) is particle-hole symmetric, i.e., our results are valid also for electron doping, $n=1+x$ (provided that one considers degenerated bands). The above states differ from those expected in the ”standard” single-band Hubbard model. Here, each holon (the site where one of the orbitals is empty) carries a $\uparrow,\downarrow$-spin flavor. Consequently, while the doping of the single-orbital Hubbard model introduces spinless holons (or doublons) in the $S=1/2$ AFM background (encapsulated in the $t$-$J$ model), for the two-orbital Hubbard model, doping introduces $S=1/2$-like objects in the $S=1$ AFM background. Note also that at the atomic level, an empty site (both orbitals empty) is possible, provided the overall desired $n$ is achieved. However, our analysis indicates that such states have vanishing weight in the ground state. In Fig. 2, we show the doping dependence of the density of various atomic-level states. We use the standard definition of (here orbital resolved) singlons and holons, i.e., $n_{\gamma\ell}^{\mathrm{S}}=n_{\gamma\ell\uparrow}+n_{\gamma\ell\downarrow}-2n_{\gamma\ell\uparrow}n_{\gamma\ell\downarrow}$ and $n_{\gamma\ell}^{\mathrm{H}}=1-n_{\gamma\ell\uparrow}-n_{\gamma\ell\downarrow}+n_{\gamma\ell\uparrow}n_{\gamma\ell\downarrow}$, respectively. Next, we define density of site singlons $n_{\mathrm{SS}}=\sum_{\ell}n_{0\ell}^{\mathrm{S}}\,n_{1\ell}^{\mathrm{S}}/L$ (sites occupied by two electrons, each at one of the orbitals), orbital singlons $n_{\mathrm{SH}}=\sum_{\ell}\left(n_{0\ell}^{\mathrm{S}}\,n_{1\ell}^{\mathrm{H}}+n_{1\ell}^{\mathrm{S}}\,n_{0\ell}^{\mathrm{H}}\right)/L$ (sites occupied by one electron at one of the orbitals), and site holons $n_{\mathrm{HH}}=\sum_{\ell}n_{0\ell}^{\mathrm{H}}\,n_{1\ell}^{\mathrm{H}}/L$. It is evident from the presented results that in the $U\gg t$ limit only site singlons and orbital singlons survive.
@@ IV Magnetic properties
As discussed in the previous section, it was shown [45, 40] that the above $S=1$ Heisenberg model description is valid not only in the $U\gg t$ limit, but also in the regions where the magnetic moments are not yet fully developed $\mathbf{S}^{2}=S(S+1)\neq 2$. For $x=0$ this is true for $U>U_{\mathrm{c}}$, while for $x\neq 0$ in the so-called orbital-RVB state, i.e., for $x\lesssim 0.3$ and $U\sim{\cal O}(W)$. Interestingly, our analysis of the average magnetic moment, $\mathbf{S}^{2}=(1/L)\sum_{i}\mathbf{S}^{2}_{i}$, indicates that in the orbital-RVB state the magnetic moments are never maximized to the $\mathbf{S}^{2}_{\mathrm{max}}(x)=2\cdot(1-2x)+0.75\cdot 2x$ value. The data presented in Fig. 3 indicates that $\mathbf{S}^{2}=\mathbf{S}^{2}_{\mathrm{max}}(x)$ only for $x=0\,,U\gg t$ and when the system orders ferromagnetically for $x\neq 0$.
@@ IV.1 Static structure factor $S(q)$
In contrast to $x=0$ - with one sharp phase transition between the trivial (paramagnetic) state and topological AKLT-like state at $U_{\mathrm{c}}$ - the finite doping $x\neq 0$ exhibits two transitions [40]. The lower one starts at $U_{\mathrm{c}}=W/2$ at $x=0$ and slowly increases till $U_{\mathrm{c}}(x)\sim 4W$ at $x\simeq 0.35$. At large $U$ (and consequently large $J_{\mathrm{H}}$), the system undergoes a second, upper transition, entering the ferromagnetic state with trivial topological properties. Our analysis of static SSF indicates that the latter transition is marked by a continuous change of the position maximum of $S(q)$ from $q_{\mathrm{max}}=2k_{\mathrm{F}}$ (incommensurate AFM) to $q_{\mathrm{max}}=0$ (ferromagnet) [see Fig. 4(b) and Fig. 4(d)]. Such behavior is akin to the spiral phase (with $\langle\mathbf{S}_{i}\times\mathbf{S}_{i+1}\rangle\neq 0$ instability) found in systems exhibiting OSMP, i.e., with lifted band degeneracy [63]. Here, for example, for $x=0.1$, we find a wide range of interaction, $3\lesssim U\lesssim 10$, where such a state is stabilized. Note that the range of interaction values that support the spiral shrinks as the doping $x$ increases [see Fig. 4(d) and sketch in Fig. 1].
@@ IV.2.1 Doping dependence
Such vanishing weight of short wavelengths $q\to\pi$ spectrum and high-$\omega$ modes were already reported in the case of the spin excitations of the OSMP systems [60, 62] (i.e., with orbital differentiation). Also, experimental inelastic neutron scattering (INS) investigation of the low-dimensional OSMP compound BaFe2Se3 [59, 65] indicates on the presence of such modes. There, the high-frequency modes take the form of well-defined (”coherent”), momentum-independent excitations. It was shown [62] that the latter depends on the value of the Hund exchange $J_{\mathrm{H}}$. Here, in the case of the doped $S=1$ Heisenberg model, the high-frequency mode is incoherent and dispersive. Nevertheless, the large frequency span $\sim 20J\sim J_{\mathrm{H}}$ could indicate that the Hund exchange also controls these high-$\omega$ features of the spectra.
IV.2.2 Hubbard interaction dependence